OpenStudy (anonymous):

What is the percent decrease from 350 to 70?

OpenStudy (tkhunny):

What is the amount decrease? That's where you need to start. Show your work.

OpenStudy (anonymous):

i first started with this 350-70=280

OpenStudy (anonymous):

and then I divided 70/280=4

OpenStudy (mimi_x3):

Great going .... I like how this is proceeding

OpenStudy (anonymous):

thanks

OpenStudy (tkhunny):

Sorry, it is a little distracting, sometimes, responding to creepers and other inappropriate people and getting them out of OpenStudy. % Change requires a BASE. Start at the base and use the base for EVERYTHING. What is the AMOUNT of change: 350 - 70 = 280 See how that started at the BASE? What is the percent change? 280 / 350 = ?? See how that is compared to the BASE?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

And the answer is....???

OpenStudy (anonymous):

i got 1.25

OpenStudy (anonymous):

please tell me if I am right

OpenStudy (tkhunny):

?? How did you get that? Did you evaluate 280/350 and then make it a percent?

OpenStudy (anonymous):

oh i divided

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

ok is it 4/5 ok i am getting confused

OpenStudy (anonymous):

how do you do this

OpenStudy (anonymous):

is it 125%

OpenStudy (anonymous):

80%

OpenStudy (anonymous):

You should only make comparisons with the original value, which is 350

OpenStudy (anonymous):

So, 350-70 shows how much change occurs. Then you compare this change to the original value, so: (350-70)/350 = 0.8 = 80%

OpenStudy (anonymous):

um i already did the test sorry